OpenStudy (anonymous):

Use the method of Lagrange multipliers to find the location (x, y) of the point on the curve y^2-x+2= that is closest to the origin.

OpenStudy (anonymous):

y^2-x+2=0 ?

OpenStudy (anonymous):

yeah =0

OpenStudy (anonymous):

distance^2= x^2+y^2 F= x^2 +y^2 grad F= {2x,2y} grad g={-1,2y} \[ grad F= \lambda grad G\]

OpenStudy (anonymous):

\[2x=-\lambda\] \[2y=\lambda 2y\] y^2-x+2=0

OpenStudy (anonymous):

thanks dear

OpenStudy (anonymous):

is it finished?

OpenStudy (anonymous):

you have to solve for x and y

OpenStudy (anonymous):

oh kul,thanks

OpenStudy (anonymous):

2y=λ2y can mean two thing 1) λ=1 2) y=0 let's assume second case y=0 y^2-x+2=0 0^2- x+2 x=2 (2,0)

OpenStudy (jamesj):

This also makes a lot of intuitive sense. Draw the graph of y^2 = x - 2 and you'll see.